a compound was found to contain 5.37% nitrogen. what is the minimum molecular mass of compound?

Grade:11

2 Answers

Chetan Mandayam Nayakar
12 years ago

PLEASE REMOVE MY PREVIOUS SOLUTION:

13n/M=0.0537

M=13n/0.0537=2905

giridharan kannan
5 years ago
According to the question,
100g of Nitrogen compound contains 5.37g of Nitrogen,
Then, for 14g of Nitrogen (molar mass of Nitrogen), the molecular mass of the compound = (100/5.37)×14=260.7g.                        (By unitary method)